OpenStudy (anonymous):

sec h + tan h. I need to express in terms of sin h and cos hand simplify. any help?

OpenStudy (anonymous):

that is suppose to say sin h and cos h and simplify.

OpenStudy (anonymous):

sech=1/cos h (by definition) tan h = sin h/ cos h (by definition) so, sec h +tan h = 1/cos h + sin h / cos h = (common denominator) = (1+sin h) /cos h